The U.S. Treasury bill has a return of 4.75 percent while the S&P 500 is returning 12.8 percent.

Your portfolio has an actual raw return of 15.6 percent and a beta of 1.43.

What is Jensen's alpha of your profile?
Answer #1

using Capital Asset Pricing Model

Required return = Rf + b ( Rm – Rf )

Where,

Rf – Risk free return = 4.75%

b – Beta = 1.43

Rm – Expected return on market portfolio = 12.8%

Required return = 4.75+1.43*(12.8-4.75)

= 4.75+11.5115

= 16.2615%

Jensen's alpha = Actual Return - Required return

= 15.6-16.2615

= -0.6615

The stock is giving a lower return than its required return, so it is advised not to accept.